OpenStudy (anonymous):

Just a quick question, can you vide zero? Some say technically you can, others say you can't.

OpenStudy (anonymous):

divide*

OpenStudy (abb0t):

wether it's infinite or undefined depends on what is being asked in the question.

OpenStudy (abb0t):

It might be reasonable to consider \(\frac{1}{0}\) = \(\infty\). But if you draw a graph of \(\frac{1}{x}\), you will see the graph shoot up to infinity just to the right of x = 0. However, just to the left, you see it shooting down to negative infinity. It could be either or and so so it;s undefined.